stone is thrown upwards with velocity v. At half the height, its velocity is
(a) zero
(b) v/4
(c) v/2
(d) v/1.414

Grade:11

1 Answers

Vikas TU
7 years ago
Total height it traversed =>
0^2 = v^2 – 2gH
H = v^2/2g...............................(1)
At half the height.
v’^2 = v^2 – 2g(H/2)
v’^2 = v^2 – v^2/2 = > v^/2 {from eqn. (1)}
v’ = v/root(2) m/s
d option.